OpenStudy (arindameducationusc):

why is h=6.63*10^-34 Js? Any derivation?

OpenStudy (aaronq):

it's derived from empirical evidence when the linear and proportional relationship between the energy in electromagnetic radiation and frequency was measured - with the proportionality constant equal to 6.626*10^-34 Js http://physics.info/planck/

OpenStudy (michele_laino):

The first value for constant h, were obtained by the same M. Planck, using the available experimental data. The value obtained by M. Planck, in the far 1901, was: \[\Large h = 6.55 \times {10^{ - 27}}erg \cdot \sec \] here is the source: \[\large \begin{gathered} {\text{M}}{\text{. Planck}}{\text{, Annalen der Physik (Lipsia) }}{\mathbf{4}}{\text{, 564 }}\left( {{\text{1901}}} \right); \hfill \\ Abhandlungen,{\text{ Vol}}{\text{. 1}}{\text{, pag}}{\text{. 728}}{\text{.}} \hfill \\ \end{gathered} \]

OpenStudy (arindameducationusc):

@Michele_Laino @aaronq Thank you for the info, I will look into that. But can you guys give some video link to understand nicely

OpenStudy (arindameducationusc):

@Robert136 need video link for this

OpenStudy (abhisar):

This one is detailed derivation of plank's constant. https://www.youtube.com/watch?v=syQbWP-7WC4

OpenStudy (arindameducationusc):

Will see that...

OpenStudy (arindameducationusc):

@Abhisar. I want why h=6.63*10^-34 I don't want Planks law! See on that if you can

OpenStudy (arindameducationusc):

Actually I got it why h=6.64*10^-34 . Anyway thanks for trying....

OpenStudy (arindameducationusc):

That's a very beautiful website which you gave @aaronq